Turing machines, closure properties, decidability Let P(z) = coz + ca-! + . . . +4-12+ cn be a polynomial with a root at r
Answer #1

(a) If x_0=0 then the inequality is immediate. Thus, we may assume that x_0\neq 0; since x_0 is an integer, this implies |x_0|\geq 1.

Since P(x_0)=0 we get

\begin{align*}0&=c_0x_0^n+c_1x_0^{n-1}+\cdots+c_{n-1}x_0+c_n\\ \Rightarrow\hspace{1.5cm}c_0x_0^n&=-c_1x_0^{n-1}-\cdots-c_{n-1}x_0-c_n\\ \Rightarrow\hspace{1.55cm}c_0x_0&=-c_1-{\frac{c_2}{x_0}}\cdots-{\frac{c_{n-1}}{x_0^{n-2}}}-{\frac{c_n}{x_0^{n-1}}}\end{align*}

Therefore,

\begin{align*}|c_0x_0|&=\begin{vmatrix}-c_1-{\frac{c_2}{x_0}}\cdots-{\frac{c_{n-1}}{x_0^{n-2}}}-{\frac{c_n}{x_0^{n-1}}}\end{vmatrix}\\ &\leq \begin{vmatrix}-c_1\end{vmatrix}+\begin{vmatrix}-{\frac{c_2}{x_0}}\end{vmatrix}+\cdots+\begin{vmatrix}-{\frac{c_{n-1}}{x_0^{n-2}}}\end{vmatrix}+\begin{vmatrix}-{\frac{c_n}{x_0^{n-1}}}\end{vmatrix}\\ &=\begin{vmatrix}c_1\end{vmatrix}+\begin{vmatrix}{\frac{c_2}{x_0}}\end{vmatrix}+\cdots+\begin{vmatrix}{\frac{c_{n-1}}{x_0^{n-2}}}\end{vmatrix}+\begin{vmatrix}{\frac{c_n}{x_0^{n-1}}}\end{vmatrix}\\ &\leq \begin{vmatrix}c_1\end{vmatrix}+\begin{vmatrix}c_2\end{vmatrix}+\cdots+\begin{vmatrix}c_{n-1}\end{vmatrix}+\begin{vmatrix}c_n\end{vmatrix}\\ &\leq n|c_{\max}|\end{align*}

which proves

\begin{align*}|x_0|&\leq {\frac{n|c_{\max}|}{|c_0|}}\end{align*}

b) By part a), to decide whether an integer polynomial \begin{align*}P(x)\end{align*} of degree \begin{align*}n\end{align*} admits an integer root we only need to evaluate the polynomial at at most \begin{align*}2n|c_{\max}|+1\end{align*} integers in\begin{align*}-n|c_{\max}|,\cdots,n|c_{\max}|\end{align*}. Evaluating an integer polynomial at an integer point can be done in polynomial (in coefficients and degree) time. Therefore, the problem is polynomial time decidable.

c) By part a), to decide whether an integer polynomial \begin{align*}P(x)\end{align*} of degree \begin{align*}n\end{align*} admits a negative integer root we only need to evaluate the polynomial at at most \begin{align*}n|c_{\max}|\end{align*} integers in\begin{align*}-n|c_{\max}|,\cdots,-1\end{align*}. Evaluating an integer polynomial at an integer point can be done in polynomial (in coefficients and degree) time. Therefore, the problem is polynomial time decidable.

d) For this polynomial, \begin{align*}|c_{\max}|=500\end{align*} and \begin{align*}n=65\end{align*} , and \begin{align*}c_0=7\end{align*} . Thus, by part a), if this polynomial has an integer root then it has an integer root \begin{align*}x_0\end{align*} satisfying

\begin{align*}|x_0|\leq {\frac {65\cdot 500}{7}}\leq 4642.9\end{align*}

Therefore, the following algorithm decides whether the given polynomial admits an integer root.

Initialize \begin{align*}x_0^+=0,~x_0^-=0\end{align*} .

If \begin{align*}x_0^+,-x_0^-<4642\end{align*}

---- evaluate \begin{align*}P(x_0^+),~P(x_0^-)\end{align*}

   ---- if \begin{align*}P(x_0^+)=0\end{align*} then stop and return polynomial has +ve integer root \begin{align*}x_0^+\end{align*}

   ---- else,

   ---- if \begin{align*}P(x_0^-)=0\end{align*} then stop and return polynomial has -ve integer root \begin{align*}x_0^-\end{align*}

---- else, \begin{align*}x_0^+\leftarrow 1+x_0^+\end{align*} and \begin{align*}x_0^-\leftarrow -1+x_0^+\end{align*} .

Return: polynomial has no integer root.